OpenStudy (anonymous):

Express the quadratic function in standard form. f(x) = −x2 + 10x + 1 f(x) = 2x2 − 16x + 38 f(x) = x2 + 2x – 4 f(x) = 5x2 − 10x + 1 f(x) = −x2 − 3x + 3 f(x) = 2x2 − 12x + 19

OpenStudy (anonymous):

standard form is a.x^2 + b.x + c so now just check from your questions the value of a, b and c like for the first one, f(x) = - x^2 +10.x + 1 a= -1 , b= 10 , c =1

OpenStudy (anonymous):

i have to put them in standard form. like f(x) = x^2 − 6x is f(x)=(x-3)^2-9

OpenStudy (anonymous):

oh nvm i got it. i feel so slow lol thank you for trying to help out :)

OpenStudy (anonymous):

lol